A person scores 81 on a test of verbal ability and 6.4 on a test of quantitative
ability. For the verbal ability test, the mean for people in general is 50 and the
standard deviation is 20. For the quantitative ability test, the mean for people in
general is 0 and the standard deviation is 5. Which is this person’s stronger
ability: verbal or quantitative? Explain your answer to a person who has never
had a course in statistics.

Answer by stanbon(75887)

A person scores 81 on a test of verbal ability and 6.4 on a test of quantitative
ability. For the verbal ability test, the mean for people in general is 50 and the standard deviation is 20.
------------------------------------------------------
For the quantitative ability test, the mean for people in general is 0 and the standard deviation is 5.
-------------------------------------------------------
Which is this person’s stronger ability: verbal or quantitative.
------------------------
Find the z score for verbal and for quantity.
VERBAL: z(81) = (81-50)/20 = 1.55
------------------------
QUANTITY: z(6.4) = (6.4-9)/5 = 1.28
-----------------------------------------
Stonger ability is Verbal.
==================================
Cheers,
stan H.
Reply to stanbon@comcast.net or algebra.com